I need help contructing a binomial distribution for n=4, p= 0.05
Answer by Edwin McCravy(20054)

n=4, p= 0.05

 x   p(x) = C(5,x)(.05)x(.95)5-x
-------------------------------------------------------
 0   p(0) = C(5,0)(.05)0(.95)5-0 = .774
 1   p(1) = C(5,1)(.05)1(.95)5-1 = .204
 2   p(2) = C(5,2)(.05)2(.95)5-2 = .021
 3   p(3) = C(5,3)(.05)3(.95)5-3 = .001
 4   p(4) = C(5,4)(.05)4(.95)5-4 = .000  
 5   p(5) = C(5,5)(.05)5(.95)5-5 = .000

Shorten the table to just two columns:

 x   p(x) 
---------
 0   .774
 1   .204
 2   .021
 3   .001
 4   .000  
 5   .000
---------
sum=1.000

Edwin